# Self-Explore to Avoid the Pit: Improving the Reasoning Capabilities of Language Models with Fine-grained Rewards

### v3

@@ Abstract
Training on large amounts of rationales (i.e., CoT Fine-tuning) is effective at improving the reasoning capabilities of large language models (LLMs). However, acquiring human-authored rationales or augmenting rationales from proprietary models is costly and not scalable. In this paper, we study the problem of whether LLMs could self-improve their reasoning capabilities. To this end, we propose Self-Explore, where the LLM is tasked to explore the first wrong step (i.e., the first pit) within the rationale and use such signals as fine-grained rewards for further improvement. On the GSM8K and MATH test set, Self-Explore achieves 11.57% and 2.89% improvement on average across three LLMs compared to supervised fine-tuning (SFT). Our code is available here.
@@ 1 Introduction
Recent works have shown that large language models (LLMs) can solve complex reasoning tasks with Chain-of-Thought (CoT) Prompting, which involves generating a rationale before its final prediction (Wei et al., 2023; Kojima et al., 2023; OpenAI et al., 2023; Team et al., 2023). In contrast, relatively smaller models show limited performance improvements, and thus prior works have focused on augmenting rationales from proprietary LLMs and distilling them to smaller models (Li et al., 2022; Fu et al., 2023; Kim et al., 2023b; Mukherjee et al., 2023; Yu et al., 2023b; Liu et al., 2023a; Mitra et al., 2023, 2024a; Li et al., 2024).
However, acquiring high-quality rationales remains challenging. For humans, hand-crafting detailed step-by-step rationale annotations is time-consuming and costly (Kim et al., 2023a). On the other hand, using close-sourced models through APIs incurs high expenses and distillation-based methods are inherently limited by the performance of their teacher model, which acts as the upper bound (Gudibande et al., 2023; Ye et al., 2023). Hence, such strategies are limited in advancing frontier models (Stanton et al., 2021; Gudibande et al., 2023). One potential solution to address this issue is to enhance the reasoning capabilities of LLMs through self-training (Gulcehre et al., 2023; Chen et al., 2024; Yuan et al., 2024).
@@
Inspired by prior works that focus on aligning LLMs to user preferences through self-training, we propose Self-Explore, a training method designed to self-improve the reasoning capabilities of LLMs by extracting granular learning signals from its own generated rationales. Specifically, a target model conducts step-level exploration to identify the first wrong step (i.e., first pit) within each rationale by sampling multiple continuations. Then, we construct a pair-wise dataset by sorting the rationales into positive and negative samples at a step level. Finally, by applying an arbitrary preference learning objective (e.g., Direct Preference Optimization (DPO) (Rafailov et al., 2023)) on a step-level, we increase the probability of generating positive rationales and lower the probability of generating negative ones in a fine-grained manner.
@@ 2.3 Self-Training for Mathematical Reasoning
Another line of works focus on self-training methods that compensate for the scarcity of high-quality training data. This includes utilizing self-generated correct rationales for training  (Zelikman et al., 2022; Yuan et al., 2023; Ni et al., 2023), and also self-generated incorrect rationales (Havrilla et al., 2024b; Hosseini et al., 2024) - which can together form a pairwise dataset that can be trained with preference learning techniques, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023).
@@ 3.1 Rejection Sampling Fine-Tuning
With the resulting model $M_{\mathrm{SFT}}$, we sample N candidate rationales $\hat{y_{i}}$ for each question with a nonzero temperature $T$ to form $\mathcal{D}_{\mathrm{GEN}}=\{(x_{i},\hat{y}_{i,j})_{j=1}^{N}\;|\;x_{i}\in Q\}$. After removing duplicate rationales using heuristics, each solution $\hat{y_{i,j}}$ is labeled as correct or incorrect by extracting their predicted final answer with extractor function $\mathcal{F}$ and comparing to the actual answer $a_{i}$. This set of correct rationales forms $\mathcal{D}_{\mathrm{RFT}}$, and $M_{\mathrm{PT}}$ is trained on this dataset with objective in eq. 1. Note that $\mathcal{D}_{\mathrm{RFT}}$ does not include initial dataset $\mathcal{D}$ in our setting.
@@ 3.2 Direct Preference Optimization
DPO (Rafailov et al., 2023) training requires a pairwise dataset consisting of a chosen completion $y^{+}$ and a rejected completion $y^{-}$ for a given input $x$. Its objective effectively increases the log-likelihood of the chosen completion over the rejected one:
@@ 4.2 Multi-Step Preference Learning
Consider a multi-step problem setting given agent $\pi_{\theta}$, input $x$, and target sequence $y$ comprised of steps $\{y^{1},...,y^{n}\}$. Then, we can naively define the agent reward by evaluating the predicted final answer at the terminal state $y^{n}$ against the ground truth answer $a$, using extractor function $\mathcal{F}$.
@@
If the answer space is large enough, we can safely assume that a match between these two indicates that all the prior steps are also correct. On the other hand, if the terminal state $y_{n}$ reached an incorrect answer, this suggests that the sequence generation has encountered at least one "pit" - an irreversible error in its prior steps $\{y^{1},...y^{i-1}\}$ that caused the agent to deviate from the correct path. Yet once we identify the first pit, we can consider all subsequent steps as non-relevant, given that they are already compromised by the preceding pit. Then, we can re-design the reward for generating each step $y^{i}$ in the multi-step problem setting as follows:
@@
Meanwhile, the challenge posed by multi-step tasks is that it is hard to avoid the pit and reach the right terminal state, especially when the problem requires many steps to solve. For simplicity, if we assume there is a constant probability of $\epsilon$ to fall into the pit in each stage, then the expected reward after generating t steps becomes $(1-\epsilon)^{t}$, which exponentially decreases as $t$ gets larger. In order to minimize this risk, previous works have utilized DPO to enable the original model as a reward model, steering away from the episodes that the model fell into the pit. However, DPO objective shown in eq. 2 relatively decreases the likelihood of all tokens in the rejected solution $y^{-}$. In light of eq. 4, we claim that only the step corresponding to the first pit should be discouraged. To elucidate, we consider the following two cases.
@@ 4.3 Self-Explore
To form a $\mathcal{D}_{\mathrm{g}\text{-}\mathrm{pair}}$ instance, we set the first pit $s^{w}$ as the new rejected sample. The new input is then created by concatenating the original input (question) with all steps prior to the first pit. For the new chosen sample, we randomly select one correct completion from $s^{w-1}$, which matches this new input. We intentionally use the whole completion from the explorer to maximize the expected reward, or the likelihood of deriving the correct answer. In a similar manner, if $w=1$, we simply use the original chosen sample. Then, we apply preference learning objective in eq. 2 on our reference model $M_{\mathrm{RFT}}$ using the new dataset $\mathcal{D}_{\mathrm{g}\text{-}\mathrm{pair}}$.
@@ 7 Conclusion
In this paper, we propose Self-Explore where LLMs can self-improve from given initial human curated dataset using fine-grained supervision. By utilizing automatic self-exploratory annotation, Self-Explore effectively integrates the roles of the annotator, target, and reward models into a single system. On mathematical reasoning datasets GSM8K and MATH, our method outperforms traditional supervised fine-tuning (SFT) method by 11.57% and 2.89% in average across three different models, respectively. We hope our work could motivate future works to explore self-training methods that could more robustly generalize to a broader reasoning space across various domains, with ends of advancing the frontier of LLM reasoning.
Limitations
We propose a method on how to better exploit the solution space to provide a better fine-grained supervision for self-improving reasoning capabilities. Yet given limited amount of questions, which is a quite common scenario, preference learning with self-generated samples may be prone to exploitation (or overfitting) and thus increases top-1 performance at the expense of diminished test-time exploration robustness. (See Appendix A). We leave as a future work to explore the potential of integrating collection of diverse datasets as in Longpre et al. (2023) to mitigate this issue, thereby enabling the model to generalize across a broader question space.
@@
Also, our work is currently conducted with 7B pre-trained models and does not consider extensively fine-tuned CoT models or larger scale architectures that have shown stronger reasoning capabilities (Yu et al., 2023b; Mitra et al., 2024b; Shao et al., 2024). We believe for practical self-training applications, it is crucial to explore continuous training processes on these sophisticated models. We encourage future works to investigate about how these open-sourced, advanced reasoning models can further enhance their self-improvement processes in a robust and effective manner.

### v4

# Self-Explore: Enhancing Mathematical Reasoning in Language Models with Fine-grained Rewards

@@ Abstract
Training on large amounts of rationales (i.e., CoT Fine-tuning) has been found effective for improving mathematical reasoning of large language models (LLMs). However, acquiring human-authored solutions or augmenting rationales from proprietary models is costly and not scalable. In this paper, we study the problem of whether LLMs could self-improve mathematical reasoning capabilities. To this end, we propose Self-Explore, where the LLM is tasked to explore the first wrong step (i.e., the first pit) within the rationale and use such signals as fine-grained rewards for further improvement. On the GSM8K and MATH test set, Self-Explore achieves 11.57% and 2.89% improvement on average across three LLMs compared to supervised fine-tuning (SFT). Our code is available here.
@@ 1 Introduction
Recent works have shown that large language models (LLMs) can solve complex reasoning tasks with Chain-of-Thought (CoT) Prompting, which involves generating a rationale before its final prediction (Wei et al., 2023; Kojima et al., 2023; OpenAI et al., 2023; Team et al., 2023). Such ability is especially evident for mathematical reasoning, where many times precise reasoning over multiple steps is required to reach the correct answer  (Fu et al., 2023; Chen et al., 2023). Meanwhile, relatively smaller models have shown limited performance, and thus prior works have focused on augmenting rationales from proprietary LLMs and distilling them to smaller models (Li et al., 2022; Kim et al., 2023b; Mukherjee et al., 2023; Yu et al., 2023b; Liu et al., 2023a; Mitra et al., 2023, 2024a; Li et al., 2024).
However, acquiring high-quality solutions remains challenging. For humans, hand-crafting detailed step-by-step rationale annotations is time-consuming and costly (Kim et al., 2023a). On the other hand, using close-sourced models through APIs incurs high expenses and distillation-based methods are inherently limited by the performance of their teacher model, which acts as the upper bound (Gudibande et al., 2023; Ye et al., 2023). Hence, such strategies are limited in advancing frontier models (Stanton et al., 2021; Gudibande et al., 2023). One potential solution to address this issue is to improve general capability of LLMs through self-training (Gulcehre et al., 2023; Chen et al., 2024; Yuan et al., 2024).
@@
Inspired by prior works that focus on aligning LLMs to user preferences through self-training, we propose Self-Explore, a training method designed to self-improve the mathematical reasoning capabilities of LLMs by extracting granular learning signals from its own generated rationales. Specifically, a target model conducts step-level exploration to identify the first wrong step (i.e., first pit) within each rationale by sampling multiple continuations. Then, we construct a pair-wise dataset by sorting the rationales into positive and negative samples at a step level. Finally, by applying an arbitrary preference learning objective (e.g., Direct Preference Optimization (DPO) (Rafailov et al., 2023)) on a step-level, we relatively increase the probability of generating positive rationales and lower the probability of generating negative ones in a fine-grained manner.
@@ 2.3 Self-Training for Mathematical Reasoning
Another line of works focus on self-training methods that compensate for the scarcity of high-quality training data. This includes utilizing self-generated correct rationales for training  (Zelikman et al., 2022; Huang et al., 2022; Yuan et al., 2023; Ni et al., 2023), and also self-generated incorrect rationales (Havrilla et al., 2024b; Hosseini et al., 2024) - which can together form a pairwise dataset that can be trained with preference learning techniques, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023).
@@ 3.1 Rejection Sampling Fine-Tuning
With the resulting model $M_{\mathrm{SFT}}$, we sample N candidate rationales $\hat{y_{i}}$ for each question with a nonzero temperature $T$ to form $\mathcal{D}_{\mathrm{GEN}}=\{(x_{i},\hat{y}_{i,j})_{j=1}^{N}\;|\;x_{i}\in Q\}$. After removing duplicate rationales using heuristics, each solution $\hat{y_{i,j}}$ is labeled as correct or incorrect by extracting their predicted final answer with extractor function $\mathcal{F}$ and comparing to the actual answer $a_{i}$. This set of correct rationales forms $\mathcal{D}_{\mathrm{RFT}}$, and $M_{\mathrm{PT}}$ is trained on this dataset with objective in eq. 1. We highlight that in domains with a sufficiently large answer space, (i.e. numeric), a correct final answer strongly indicates that the rationale is likely error-free, which is a notable advantage of mathematical reasoning tasks. [^1] In contrast, if the answer space is small (i.e. true/false or multiple choice) selecting the correct option does not necessarily guarantee that the rationale is also correct.
@@ 3.2 Direct Preference Optimization
DPO (Rafailov et al., 2023) training requires a pairwise dataset consisting of a chosen completion $y^{+}$ and a rejected completion $y^{-}$ for a given input $x$. Its objective relatively increases the log-likelihood of the chosen completion over the rejected one:
@@ 4.2 Multi-Step Preference Learning
In preference learning, a language model $\pi_{\theta}$ functions as an agent optimized to generate responses that maximize a given reward function. In a multi-step problem setting, given an input $x$ and a target sequence $y$ comprising steps $\{y^{1},...,y^{n}\}$, we can define the agent’s reward by evaluating the predicted final answer at the terminal state $y^{n}$ against the ground truth answer $a$ using the extractor function $\mathcal{F}$
@@
If the answer space is large enough, we can safely assume that a match between these two indicates that all the prior steps $\{y^{1},...,y^{n-1}\}$ are also correct. On the other hand, if the terminal state $y_{n}$ reached an incorrect answer, this suggests that the sequence generation has encountered at least one "pit" - an irreversible error in its prior steps that caused the agent to deviate from the correct path. Yet once we identify the first pit, we can consider all subsequent steps as non-relevant, given that they are already compromised by the preceding pit. Then, we can re-design the reward for generating each step $y^{i}$ in the multi-step problem setting as follows:
@@
Meanwhile, the challenge posed by multi-step tasks is that it is hard to avoid the pit and reach the correct terminal state, especially when the problem requires many steps to solve. For simplicity, if we assume there is a constant probability of $\epsilon$ to fall into the pit in each stage, then the expected reward after generating t steps becomes $(1-\epsilon)^{t}$, which exponentially decreases as $t$ gets larger. In order to minimize this risk, previous works have utilized DPO to enable the original model as a reward model, steering away from the episodes that the model fell into the pit. However, DPO objective shown in eq. 2 relatively decreases the likelihood of all tokens in the rejected solution $y^{-}$. In light of eq. 4, we claim that only the step corresponding to the first pit should be discouraged. To elucidate, we consider the following two cases.
@@ 4.3 Self-Explore
To form a $\mathcal{D}_{\mathrm{g}\text{-}\mathrm{pair}}$ instance, we set the first pit $s^{w}$ as the new rejected sample. The new input is then created by concatenating the original input (question) with all steps prior to the first pit. For the new chosen sample, we randomly select one correct completion from the step just before the first pit ($s^{w-1}$), that matches this new input. We intentionally use the whole completion from the explorer to maximize the expected learning signal, thus the likelihood of deriving the correct answer. In a similar manner, if $w=1$, we simply use the original chosen sample. Finally, we train with preference learning objective in eq. 2 on our reference model $M_{\mathrm{RFT}}$ using this new dataset $\mathcal{D}_{\mathrm{g}\text{-}\mathrm{pair}}$.
@@ 7 Conclusion
In this paper, we propose Self-Explore where LLMs can self-improve from given initial human curated dataset using fine-grained supervision. By utilizing automatic self-exploratory annotation, Self-Explore effectively integrates the roles of the annotator, target, and reward models into a single system. On mathematical reasoning datasets GSM8K and MATH, our method outperforms traditional supervised fine-tuning (SFT) method by 11.57% and 2.89% in average across three different models, respectively. Furthermore, we demonstrate that our method introduces minimal computation overhead (See Appendix H). We hope our work could motivate future works to explore self-training methods that could more robustly generalize to a broader reasoning space across various domains, with ends of advancing the frontier of LLM reasoning.
@@
Limitations
We propose a method on how to better exploit the solution space to provide a better fine-grained supervision for self-improving reasoning capabilities. Yet given limited amount of questions, which is a quite common scenario, preference learning with self-generated samples may be prone to overconfidence and thus increases top-1 performance at the expense of diminished test-time exploration robustness (Cao et al., 2024). We suspect this is related to reward overoptimization (Gao et al., 2022; Burns et al., 2023) and attach relevant analysis in Appendix. We leave as a future work on methods for mitigating this overoptimization, where one promising direction could be exploring the potential of integrating collection of diverse datasets as in Longpre et al. (2023), so that the model can generalize across a broader question space.
Also, our work is currently conducted with 7B pre-trained models and does not consider extensively fine-tuned CoT models or larger scale architectures that have shown stronger reasoning capabilities (Yu et al., 2023b; Mitra et al., 2024b; Shao et al., 2024). We believe for practical self-training applications, it is crucial to explore continual training processes on these sophisticated models. While this paper aims to compensate for distilled rationales used in instruction-tuned models, we encourage future works to investigate about such models could further benefit from self-improvement processes in a robust and effective manner.
Acknowledgements
This work was partly supported by the LG AI Research grant (Self-improving logical reasoning capabilities of LLMs, 2022, 50%) and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2024-00397966, Development of a Cybersecurity Specialized RAG-based sLLM Model for Suppressing Gen-AI Malfunctions and Construction of a Publicly Demonstration Platform, 50%).
@@ Appendix G FLASK Prompt
Appendix H Computational Costs
We report the overall computational costs (i.e. GPU Hours) for each baseline, including the exploration stage in Table 8. Our baselines involve different training (Tr.) and generation (Gen.) stages. Note that the table reflects the following configuration:
  • Mistral 7B
  • 5 epoch SFT & RFT training
  • 3 epoch DPO training
  • 7.4M RFT samples generated
  • 39K pairwise-data exploration
@@ Appendix H Computational Costs
Stages SFT (Tr.) RFT (Gen.) RFT (Tr.) Exploration (Gen.) DPO (Tr.) Total Time (hr)
GPU Hours 1.3 6 11.2 2.7 20
SFT 1.3
RFT 18.5
DPO 38.5
Self-Explore 41.2
Table 8: GPU hours for different baselines.